# Exploring Homogeneous and Heterogeneous Consistent Label Associations for Unsupervised Visible-Infrared Person ReID

### v2

@@ 4 Experiments
Table 3: Comparison with other cross-modality label association methods for unsupervised VI-ReID on SYSU-MM01. All experiments are based on our MIRL framework (Sec.3.2).
@@
Method All-search Indoor-search R1 mAP mINP R1 mAP Baseline 41.76 38.57 24.32 45.83 54.43 MIRL (Sec.3.2) w/o OCLR (Sec.3.3) w/ PGM [46] 48.30 43.09 27.67 45.65 53.73 w/ BCCM [5] 53.80 47.55 30.53 50.72 59.41 w/ DOTLA [6] 59.27 50.60 32.36 55.34 62.76 w/ MULT (Ours) 61.48 54.43 38.02 60.73 66.32 MIRL (Sec.3.2) w/ OCLR (Sec.3.3) w/ PGM [46] 54.54 49.88 34.01 54.94 62.98 w/ BCCM [5] 55.64 50.69 34.79 57.80 65.47 w/ DOTLA [6] 61.74 54.65 38.61 59.19 66.84 w/ MULT (Ours) 65.03 58.62 42.77 65.35 71.24
4.1 Dataset and Evaluation Protocol
Datasets. Our proposed method is evaluated on two public visible-infrared datasets, namely SYSU-MM01 [44] and RegDB [33]. SYSU-MM01 comprises 395 identities, with 22258 visible and 11909 infrared images captured by indoor and outdoor cameras. RegDB is a smaller dataset obtained from a pair of aligned visible and infrared cameras [33]. It contains 412 identities, where each identity has 10 visible images and 10 infrared images.
Evaluation Metrics. All experiments follow the common evaluation protocols used for VI-ReID [51, 49, 50]. The evaluation metrics include Cumulative Matching Characteristic (CMC), Mean Average Precision (mAP), and Mean Inverse Negative Penalty (mINP [51]). For the SYSU-MM01 dataset, we evaluate the proposed method under two search modes: the All-search mode and the Indoor-search mode. As for the RegDB dataset, we evaluate our method on two testing modes: Visible-to-Infrared and Infrared-to-Visible. We randomly selected 206 identities for training and the remaining 206 for testing.
Implementation Details. Our proposed method is implemented using PyTorch. we adopt two-stream ResNet50 [17] pretrained on ImageNet [12] as our backbone. In a mini-batch, the number of classes $P$ and instances for each class $K$ are both 12. All the input images are resized to $288\times 144$. The augmentations for images are following [47]. Besides, Linear Transform Generator (LTG [38]) and ColorJitter augmentations are adopted in the DCL training stage. We train for a total of 90 epochs. The DCL framework is trained in the first 40 epochs, while our proposed framework is trained in the other 50 epochs. We use the Adam optimizer for training the model with weight decay 5e-4. The initial learning rate is set to 3.5e-4 and decays 10 times at the 20th, 60th, and 80th epochs. Following [5, 47, 46], the momentum factor $\mu$ is 0.1, and the temperature $\tau$ is 0.05, while the maximum distance for DBSCAN is set to 0.6 on SYSU-MM01 and 0.3 on RegDB. The parameter $\kappa$ for $\kappa$-reciprocal nearest neighbors in Eq.5 is set to 30 following [57]. The hyper-parameter $\lambda$ in Eq.6 for OT is set to 25 following [42]. The trade-off parameter $\alpha$ in Eq.10 is set to 0.2 and $\beta$ in Eq.13 is set to 0.7.
@@ 4.1 Dataset and Evaluation Protocol
4.2 Comparison with State-of-the-Arts
To demonstrate the efficiency of our proposed methods, we compare our method with state-of-the-art methods under three relevant settings on SYSU-MM01 and RegDB datasets, i.e., supervised VI-ReID methods, unsupervised single-modality ReID methods, and unsupervised VI-ReID methods. The results are shown in Tab.1.
Comparison with Supervised VI-ReID Methods. Our proposed method achieves competitive performance with supervised method VCD+VML [40] on SYSU-MM01, while demonstrating excellent performance close to the state-of-the-art supervised methods on RegDB.
Comparison with Conventional Unsupervised Single-Modality ReID Methods. The results in Tab.1 indicate that unsupervised single-modality methods cannot effectively address the large modality-gap in cross-modality scenarios, emphasizing the necessity of proposing cross-modality label association strategies for VI-ReID.
Comparison with Unsuervised VI-ReID Methods. As reported in Tab.1, our method outperform the state-of-the-art GRU [48] by 1.63% mAP, 4.08% Rank-1 on SYSU-MM01 (All-search) and 13.5% mAP, 17.59% Rank-1 on RegDB (Visible-to-infrared). The proposed method has three main advantages: (1) Unlike H2H [24] and OTLA [42], our method does not require additional datasets or camera labels for training. (2) Our MULT associates cross-modality labels from a novel perspective and can be applied to other unsupervised cross-domain tasks. (3) Our OCLR is a plug-and-play module that can be employed in other unsupervised memory-based methods to handle label noise.
@@ 4.2 Comparison with State-of-the-Arts
4.3 Ablation Study
To validate the effectiveness of each component of our method, we conduct ablation experiments on SYSU-MM01 and RegDB datasets, as shown in Tab.2. Our baseline consists of a two-stream ResNet trained under the DCL framework [47]. All experiments are conducted with our MULT module. We did not perform an ablation study on $\mathcal{L}_{IM}$ as it does not address the problem of modality discrepancy.
Effectiveness of MULT. The MULT module delivers a +15.41% mAP and +18.67% Rank-1 improvement by directly incorporating $\mathcal{L}_{CM}$ (see $1^{st}$ row and $2^{nd}$ row in Tab.2) and achieves +20.05% mAP and +23.27% Rank-1 improvement when trained within our entire framework on SYSU-MM01 (see $1^{st}$ row and $7^{th}$ row in Tab.2). To further evaluate the effectiveness of MULT, we compare MULT with other cross-modality label association methods in Tab.3, including PGM [46], MBCCM [5] and DOTLA [6]. The results demonstrate our MULT provides higher-quality associations than other methods.
Effectiveness of OCLR. The OCLR module provides a performance gain of +8.8% mAP and +10.12% Rank-1 when directly adding it to our baseline (see $1^{st}$ row and $4^{th}$ row in Tab.2). When used in conjunction with our MIRL framework, the OCLR further improves the performance of +4.19% mAP and +3.55% Rank-1 (see $3^{rd}$ row and $7^{th}$ row in Tab.2). OCLR utilizes ever-evolving prototypes for online refinement, avoiding overfitting noisy labels. We also integrate our OCLR into other methods (see upper and lower in Tab.3) based on our MIRL framework. The results indicate that our OCLR improves performance when collaborating with other cross-modality association methods.
Effectiveness of MIRL. Our MIRL achieves +0.45% mAP improvement without OCLR (see $2^{nd}$ row and $3^{rd}$ row in Tab.2) and achieves +2.41% mAP with OCLR (see $6^{th}$ row and $7^{th}$ row in Tab.2) compared to leveraging $\mathcal{L}_{CM}$ alone. The results illustrate the importance of incorporating intra-modality contrastive loss $\mathcal{L}_{IM}$ into our framework.
4.4 Further Analysis
Parameter Analysis. The proposed method includes two parameters in MULT, i.e., $\alpha$ in Eq.10 and $\beta$ in Eq.13. To investigate the impact of these two parameters, we varied their values, as depicted in Figure 3. We find when $0.15\leq\alpha\leq 0.3$ and $0.6\leq\beta\leq 0.7$, the model achieves relatively outstanding performance. We set $\alpha=0.2$ and $\beta=0.7$ based on the experiments.
Pseudo-label Accuracy Analysis. To quantify the quality of the pseudo-labels, we calculate the accuracy of positive instance pairs found by the pseudo-labels during training, as shown in Fig.4. We compare our MULT with DOTLA [6], and the results demonstrate that our MULT significantly facilitates the quality of pseudo labels. Simultaneously, the results indicate that as MULT and network training alternate, the network can continuously correct erroneous pseudo-labels.
Visualization. We visualize the Euclidean distance distribution of randomly selected 50000 positive and negative visible-infrared pairs, as shown in Fig.5. By sequentially integrating the modules (i.e., MULT, MIRL, and OCLR) into the training framework, we observe a convergence of cross-modality positive pairs and a divergence of negative pairs, demonstrating the effectiveness of each component in our framework to address the modality discrepancy. Fig.5 (c) and Fig.5 (d) further illustrate the superiority of MULT in comparison with DOTLA [6].
@@ 4.4 Further Analysis
Figure 5: The visualization of distance distribution of randomly selected cross-modality positive and negative pairs.
@@ 5 Conclusion
In this paper, we propose the Modality-Unified Label Transfer module to establish high-quality cross-modality pseudo-label associations for training. Our MULT lays emphasis on both homogeneous and heterogeneous context structure consistency by instance-wise affinities-guided label transfer. Furthermore, we introduce a straightforward yet effective plug-and-play Online Cross-memory Label Refinement module, which simultaneously mitigates the impact of noisy labels and facilitates modality alignment. Extensive experiments have demonstrated that our framework outperforms prior state-of-the-art methods.
Limitation. One major limitation of this study is there are still inevitable noisy labels from MULT and there is considerable potential for improvement in comparison to the supervised VI-ReID methods. In the future, we will investigate more robust cross-modality label association methods.

### v4

@@ 4.3 Ablation Study
Table 4: Comparison with other cross-modality label association methods for unsupervised VI-ReID on SYSU-MM01. All experiments are based on our AMIRL framework (Sec.3.2).
@@
Method Venue All-search Indoor-search R1 mAP mINP R1 mAP mINP Baseline - 41.76 38.57 24.32 45.83 54.43 50.00 AMIRL (Sec.3.2) w/o OCLR (Sec.3.3) w/ PGM Wu and Ye (2023) CVPR-23 50.85 45.50 30.64 51.31 59.73 55.34 w/ BCCM Cheng et al (2023a) MM-23 51.35 46.61 31.97 53.44 61.28 56.66 w/ CLU Yang et al (2023) ICCV-23 57.09 48.96 31.88 57.11 64.51 60.32 w/ DOTLA Cheng et al (2023b) MM-23 57.17 51.33 36.07 55.16 62.53 57.87 w/ MULT (Ours) - 60.18 55.52 39.86 60.48 67.39 62.94 AMIRL (Sec.3.2) w/ OCLR (Sec.3.3) w/ PGM Wu and Ye (2023) CVPR-23 56.43 52.31 37.00 59.28 66.04 61.55 w/ BCCM Cheng et al (2023a) MM-23 58.46 53.69 37.99 58.84 66.56 62.25 w/ CLU Yang et al (2023) ICCV-23 61.40 54.18 38.11 62.91 68.32 64.08 w/ DOTLA Cheng et al (2023b) MM-23 61.53 55.35 39.96 60.82 67.60 63.06 w/ MULT (Ours) - 64.77 59.23 43.46 65.34 71.46 67.83
Table 5: Ablation study for the auxiliary memory.
Method All-search Indoor-search R1 mAP mINP R1 mAP w/ Intra 63.77 58.61 43.04 64.54 71.38 w/ Aux 63.06 58.20 42.65 65.06 71.07 Intra & Aux 64.77 59.23 43.46 65.34 71.46
Table 6: Ablation study for the alternative scheme.
Method All-search Indoor-search R1 mAP mINP R1 mAP w/ MIRL 64.05 58.44 43.06 64.86 70.90 w/ AMIRL 64.77 59.23 43.46 65.34 71.46
Effectiveness of OCLR. The OCLR module provides a performance gain of +12.46% mAP and +12.58% Rank-1 when directly adding it to our baseline (see $1^{st}$ row and $4^{th}$ row in Tab.3). When used in conjunction with our AMIRL framework, the OCLR further improves the performance of +3.71% mAP and +4.59% Rank-1 (see $3^{rd}$ row and $7^{th}$ row in Tab.3). The OCLR module significantly mitigates the impact of pseudo-label noise while narrowing the modality discrepancy. OCLR utilizes ever-evolving prototypes for online refinement, avoiding overfitting static noisy labels. We also integrate our OCLR into other methods (see upper and lower in Tab.4) based on our AMIRL framework. The results indicate that our OCLR improves performance when collaborating with other cross-modality association methods. The proposed OCLR has the following advantages compared to the unsupervised single-modality ReID method MMT Ge et al (2020a): (1) MMT focuses on consistency in predictions from different augmentations of the same instance, which does not consider the cross-modality interactions. Conversely, our OCLR emphasizes consistency in relationships between instances and multi-modality prototypes, which enhances cross-modality interactions and prompts learning modality-shared information. Therefore, the proposed OCLR is more suitable for the USL-VI-ReID task. (2) MMT refines predictions using the moving average model, while OCLR conducts refinement across multi-memories without storing past models, significantly reducing GPU memory consumption.
@@
Table 7: Comparison with various designs for modeling $\mathbf{S}^{he}$.
Method SYSU-MM01 (All-search) SYSU-MM01 (Indoor-search)
R1 mAP mINP R1 mAP mINP
Jaccrad 61.08 55.69 39.20 62.63 68.49 63.78
Cosine (k=20) 62.19 57.97 42.38 64.54 70.88 66.57
Cosine (k=30) 61.73 57.66 41.91 64.65 70.59 66.08
OT problem 64.77 59.23 43.46 66.67 71.87 67.20
Effectiveness of AMIRL. Our AMIRL achieves +0.60% mAP improvement without OCLR (see $2^{nd}$ row and $3^{rd}$ row in Tab.3) and achieves +2.41% mAP with OCLR (see $6^{th}$ row and $7^{th}$ row in Tab.3) compared to leveraging $\mathcal{L}_{CM}$ alone. The results illustrate the importance of incorporating intra-modality contrastive loss $\mathcal{L}_{IM}$ into our framework, especially when training with our OCLR module. Moreover, we conduct the ablation experiment for the auxiliary intra-modality memory and the alternative training scheme, as shown in Tab.5 and Tab.6. In Tab.5, “w/ Intra” denotes only intra-modality memory banks $\tilde{\mathbf{M}}^{v}$ and $\tilde{\mathbf{M}}^{r}$ are involved in $\mathcal{L}_{IM}$ during training, and “w/ Aux” denotes only auxiliary memory banks $\hat{\mathbf{M}}^{r}$ and $\hat{\mathbf{M}}^{v}$ are utilized. The results indicate that the cross-modality pseudo-labels also contribute to intra-modality learning. The improvement is attributed to these two types of pseudo-labels revealing two distinct intra-modality structures in feature space, thus correcting each other mutually. In Tab.6, “MIRL” denotes the contrastive learning framework without the alternative scheme (AMIRL) for V-based training and R-based training. The results demonstrate that the alternative scheme further enhances the performance. This is achieved by alternating the utilization of labels from a single-directional MULT (V2R or R2V) to supervise model training across different epochs, thereby effectively mitigating the issue of pseudo-label inconsistency between V2R and R2V MULT. Compared to a basic framework, MSMA Cheng et al (2023a), we construct an auxiliary memory bank to learn the structure of cross-modality pseudo-labels, this strategy fully exploits the pseudo-labels from MULT. Furthermore, we propose to alternate V-based training and R-based training. Such an alternative training scheme mitigates the side-effects of noisy labels from the inconsistent pseudo-labels output of the V2R / R2V MULT. our AMIRL makes more effective utilization of the cross-modality pseudo-labels $\tilde{\mathbf{y}}^{v}$ and $\tilde{\mathbf{y}}^{r}$, thus is more suitable for our MULT label association algorithm.
4.4 Further Analysis
Analysis for heterogeneous affinities modeling. The Optimal Transport (OT) problem in Eq.3 is formulated to model heterogeneous affinities $\mathbf{S}^{he}$ from the perspective of the transition between instances from two different distributions. The constraint in Eq.3 ensures that $N^{v}$ visible instances’ affinities are uniformly distributed to $N^{r}$ infrared instances in the transport plan, and vice versa. This design, compared to other alternative designs, prevents degenerated solutions where few infrared instances hold extremely high affinities with the majority of visible instances. We conducted a comparative analysis of various designs for modeling $\mathbf{S}^{he}$, as shown in Tab.7. “Jaccard” denotes the cross-modality Jaccard Similarity-based affinities and “Cosine (k=K)” denotes the cosine-similarity-based knn-affinities, K denotes the number of knn-neighbors for constructing knn-graph. The results confirmed the rationality of our choice.
Parameter Analysis. The proposed method includes two parameters in MULT, $i.e.$, $\alpha$ in Eq.14 and $\beta$ in Eq.21. $\alpha$ is a trade-off parameter between the self-consistency terms and the heterogeneous consistency terms. $\beta$ is the parameter to control the smoothness of the soft pseudo-labels from MULT. To investigate the impact of these two parameters, we varied their values, as depicted in Figure 4. We find when $0.15\leq\alpha\leq 0.3$ and $0.6\leq\beta\leq 0.7$, the model achieves relatively outstanding performance. We set $\alpha=0.2$ and $\beta=0.7$ based on the experiments.
Table 8: Comparison with other methods on FMI/ARI.
Method FMI metric (%) ARI metric (%)
RGB ($\tilde{y}^{v}$) IR ($\tilde{y}^{r}$) $\tilde{y}^{v}\&\hat{y}^{r}$ $\hat{y}^{v}\&\tilde{y}^{r}$ RGB ($\tilde{y}^{v}$) IR ($\tilde{y}^{r}$) $\tilde{y}^{v}\&\hat{y}^{r}$ $\hat{y}^{v}\&\tilde{y}^{r}$
PGM (CVPR23) 53.57 82.03 56.51 52.05 51.20 81.84 56.30 50.34
BCCM (MM23) 51.99 83.33 60.20 52.61 50.03 83.20 52.29 44.92
OTLA (ECCV22) 58.39 84.40 55.57 51.77 56.40 84.08 55.21 49.29
MULT (Ours) 65.86 86.61 69.32 62.87 63.74 86.48 69.18 61.07
Table 9: Ablation study on CA Ye et al (2021a) in the proposed OCLR on SYSU-MM01.
Method SYSU-MM01 (All-search) SYSU-MM01 (Indoor-search)
R1 mAP mINP R1 mAP mINP
Ours w/o OCLR 60.18 55.52 39.18 60.66 66.92 62.24
Ours w/ OCLR (w/o CA) 61.98 57.44 42.05 63.01 69.62 65.42
Ours w/ OCLR (w/ CA) 64.77 59.23 43.46 65.34 71.46 67.83
Table 10: Ablation study on LTG Tan et al (2023) on SYSU-MM01.
Method RegDB (Visible-to-Infrared) SYSU-MM01 (All-search)
R1 mAP mINP R1 mAP mINP
GUR(ICCV23,SOTA) 73.91 70.23 58.88 60.95 56.99 41.85
Ours w/o LTG 88.20 81.21 66.24 62.57 57.90 42.25
Ours w/ LTG 89.95 82.09 67.29 64.77 59.23 43.46
Pseudo-label Accuracy Analysis. To quantify the quality of the pseudo-labels, we calculate the accuracy/recall of positive instance pairs found by the pseudo-labels during training, as shown in Fig.5 and Fig.6. We denote the ground-truth hard labels as $y^{e}(gt)$, where $e=\{v,r\}$ indicates the visible and infrared modality, respectively. We utilize the hard form $y^{e}=\{\tilde{y}^{v},\tilde{y}^{r},\hat{y}^{v},\hat{y}^{r}\}$ of the soft pseudo-labels from MULT when quantifying the quality. In Fig.5, ‘Intra-modality visible label accuracy’ ($IntraAcc^{v}$) indicates the accuracy of the visible intra-modality positive pairs found by the cross-modality visible labels $\hat{{y}}^{v}$. This can be achieved by computing the proportion of the ground-truth positive pairs within all instance pairs that share the same pseudo-label:
$IntraAcc^{v}=\frac{\sum_{i=1}^{N^{v}}\sum_{j=1}^{N^{v}}\mathbf{1}(\hat{{y}}^{v}_{i}=\hat{{y}}^{v}_{j})\mathbf{1}({y}^{v}_{i}(gt)={y}^{v}_{j}(gt))}{\sum_{i=1}^{N^{v}}\sum_{j=1}^{N^{v}}\mathbf{1}(\hat{{y}}^{v}_{i}=\hat{{y}}^{v}_{j})}.$ (30)
Similarly, ‘Cross-modality visible label accuracy’ ($CrossAcc^{v}$) indicates the accuracy of the cross-modality positive pairs found by $\tilde{\mathbf{y}}^{v}$ and $\hat{\mathbf{y}}^{r}$ from the V2R MULT. It can be similarly formulated as follows:
$CrossAcc^{v}=\frac{\sum_{i=1}^{N^{v}}\sum_{j=1}^{N^{r}}\mathbf{1}(\tilde{{y}}^{v}_{i}=\hat{{y}}^{r}_{j})\mathbf{1}({y}^{v}_{i}(gt)={y}^{r}_{j}(gt))}{\sum_{i=1}^{N^{v}}\sum_{j=1}^{N^{r}}\mathbf{1}(\tilde{{y}}^{v}_{i}=\hat{{y}}^{r}_{j})}.$ (31)
The recall in Fig.6 can be obtained similarly by computing the proportion of pairs that share the same pseudo-label within all ground-truth positive pairs. We compare our MULT with DOTLA Cheng et al (2023b), and the results demonstrate that our MULT significantly facilitates the quality of pseudo labels. Simultaneously, the results indicate that as MULT and network training alternate, the network can continuously correct erroneous pseudo-labels.
To further measure the reliability of pseudo-labels derived from our MULT, we employ Fowlkes-Mallows Index (FMI) and Adjusted Rand Index (ARI) metrics to evaluate the pseudo-labels across various association methods, as shown in Tab.8. The FMI score indicates the geometric mean of the accuracy and recall, while the ARI score reflects the degree of overlap between the ground-truth label space and the pseudo-label space. The results illustrate the superiority of our MULT.
@@ 4.4 Further Analysis
Figure 6: Recall of intra-modality and cross-modality positive pairs found by pseudo-labels on SYSU-MM01.
@@
Figure 7: The visualization of Euclidean distance distribution of randomly selected cross-modality positive and negative pairs.
Figure 8: Visualization of the ranking lists on the SYSU-MM01 dataset. The persons who are different from the query persons are marked with red boxes, while those who are the same as the query are marked with green boxes.
Distribution Visualization. We visualize the Euclidean distance distribution of randomly selected 50000 positive and negative visible-infrared pairs, as shown in Fig.7. By sequentially integrating the modules ($i.e.$, MULT, AMIRL, and OCLR) into the training framework, we observe a convergence of cross-modality positive pairs and a divergence of negative pairs, demonstrating the effectiveness of each component in our framework to address the modality discrepancy. We further visualize the distribution by integrating DOTLA into our framework in Fig.7 (d), the results illustrate the superiority of MULT in comparison with DOTLA Cheng et al (2023b).
Figure 9: The t-SNE Van der Maaten and Hinton (2008) visualization of the feature of 7 randomly selected identities. Different colors mean different identities. Circle means the visible modality and cross means the infrared modality.
T-SNE Visualization. We visualize the t-SNE Van der Maaten and Hinton (2008) map of 7 randomly selected identities, as shown in Fig.9. In Fig.9 (a) and Fig.9 (b), it is evident that, owing to the precise label associations from MULT, cross-modality positive pairs exhibit increased proximity within the embedding space. Fig.9 (c) and Fig.9(d) provide a comparative analysis between our MULT and DOTLA Cheng et al (2023b). The results reveal the issue of homogeneous inconsistency in DOTLA, where images of the same identities are distributed into distinct clusters within the embedding space (black circles in Fig.9 (d)). Our MULT effectively addresses this problem (black circles in Fig.9 (c)) by generating pseudo-labels with low homogeneous consistency. The effectiveness of our OCLR is demonstrated when comparing Fig.9 (b) and Fig.9 (c). Without OCLR, the network struggles to learn compact and discriminative representations due to the inevitable noisy labels (red dotted circles in Fig.9 (b)). With OCLR, it shows more discriminative features that are well-distributed around the edge of the clusters (green dotted circles in Fig.9 (c)). The results indicate that our OCLR module is robust to noise while facilitating modality alignment.
Analysis of the data augmentation strategies in the proposed method. Data augmentations are widely-used in VI-ReID, we conduct ablation studies for the data augmentations included in our framework, $i.e.$, CA Ye et al (2021a) and LTG Tan et al (2023). CA Ye et al (2021a) is a commonly-used technology in existing USL-VI-ReID baselines Cheng et al (2023a); Wu and Ye (2023); Yang et al (2022); Cheng et al (2023b); Yang et al (2023). We conduct ablation studies of the CA augmentation utilized in OCLR, as shown in Tab.9. As shown in Tab.9, the OCLR module is still effective when removing CA, which outperforms the framework without OCLR by 1.92% mAP and 1.8% Rank1. When incorporating the CA-augmented images into our OCLR, the performance further improves by 1.79% mAP and 2.79% Rank1. It can be attributed to the integration of CA bridges different modalities and promotes the model to better learn modality-shared features. We further study the improvement brought by the LTG augmentation, as shown in Tab.10. The results show a slight drop in performance after removing LTG, but it still outperforms the state-of-the-art GUR. The results also show that LTG further improves the model performance.
Visualization of Ranking List. We visualize some of the ranking lists on the SYSU-MM01 dataset, as shown in Fig.8. We compare our method with the baseline pretrained under the DCL framework Yang et al (2022). The results demonstrate the effectiveness of our method in generating high-quality cross-modality supervision signals for training.
@@ 5 Conclusion
In this paper, we propose the Modality-Unified Label Transfer module to establish high-quality cross-modality pseudo-label associations for training from a novel structural consistency perspective. Our MULT lays emphasis on preserving both homogeneous and heterogeneous structure information in pseudo-label space by instance-wise affinities-guided label transfer. To fully exploit the soft pseudo-labels from MULT, an Alternative Modality-Invariant Representation Learning framework is proposed based on both intra-modality and cross-modality contrastive learning. Furthermore, we introduce a straightforward yet effective plug-and-play Online Cross-memory Label Refinement module, which simultaneously mitigates the negative effects of noisy labels and facilitates modality alignment. Extensive experiments have demonstrated that our framework outperforms prior state-of-the-art methods. In the future, we will investigate more robust cross-modality label association methods based on our proposed MULT, which is the core issue of the unsupervised VI-ReID task.
Data Availability Statement
SYSU-MM01: A signed dataset release agreement must be send to wuancong@gmail.com or wuanc@mail.sysu.edu.cn to obtain a download link.
RegDB: The dataset can be downloaded by submitting a copyright form to http://dm.dongguk.edu/link.html.